# Adopt a PET! An Exploration of PETs, Policy, and Practicalities for Industry in Canada

### v1

@@ Abstract.
Privacy enhancing technologies (PETs) are technical solutions for privacy issues that exist in our digital society. Despite increased privacy challenges and a corresponding increase in new regulations being proposed by governments across the globe, a low adoption rate of PETs persists. In this work, we investigate the relationship that new privacy regulations have on industry’s decision-making processes as well as the extent to which privacy regulations inspire the adoption of PETs. We conducted a qualitative survey study with 22 industry participants from across Canada to investigate how businesses in Canada make decisions to adopt novel technologies and how new privacy regulations impact their business processes. Through this study, we identify the breadth of approaches employed by organizations considering PETs and the challenges they face in their efforts to ensure compliance with all pertinent laws and regulations. We further identify a gap between how companies think of privacy technologies and how researchers think of privacy technologies that can contribute to low adoption of the increasingly sophisticated privacy technologies produced by researchers, such as applications of differential privacy, multiparty computation, and trusted execution environments. Informed by the results of our analysis, we make recommendations for industry, researchers, and policymakers on how to support what each of them seeks from the other when attempting to improve digital privacy protections. By advancing our understanding of what challenges industry faces in ensuring compliance with novel and existing privacy regulations, we increase the effectiveness of future privacy research that aims to help overcome these issues.
Privacy, Regulations, Industry Adoption, Privacy-Enhancing Technologies
journalyear: YYYYjournalvolume: 2025journalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Demand for privacy is influencing law-makers, resulting in ongoing introductions of new privacy regulations globally (ccpa2018, ; gdpr2016, ; billc27, ; edwards2021eu, ). Most Canadian companies are moderately aware of their responsibilities under Canada’s privacy laws and have taken steps to ensure they comply with these laws according to the most recent survey (2023-2024) of Canadian businesses (phoenixspi2024survey, ). The survey also indicates that “awareness of the OPC’s (Office of the Privacy Commissioner of Canada) information and tools for businesses has increased significantly, but only $26\%$ of the businesses have used these resources to comply with privacy obligations. The use continues to fall short of awareness due, at least in part, to lack of need”. This suggests there is a clear gap between regulatory resources and business’ utilization of them. Similarly, while there has been a lot of effort in privacy enhancing technologies (PETs) research to develop deployable privacy-enhancing tools, it is unclear whether these additional regulations will motivate increased usage of these tools by industry. Thus, we want to identify any existing or perceived impediments to adopting these technologies and complying with these regulations to better understand how the industry-research relationship can be improved.
Personal Information Protection and Electronic Documents Act (PIPEDA) (pipeda2000, ) is Canada’s federal privacy law for private-sector organizations. It sets out the ground rules for how businesses must handle personal information in the course of their commercial activity. The Canadian government proposed Bill C27 (billc27, ) in 2022 to enact the Consumer Privacy Protection Act, the Personal Information and Data Protection Tribunal Act and the Artificial Intelligence and Data Act (AIDA). While, the federal privacy law reforms and artificial intelligence regulation contained in the bill are put in an uncertain state following the prorogation of Parliament in January 2025, their potential impact on Canada’s industry cannot be ignored (iapp2025billc27, ). Further, each of Canada’s provinces and territories has their own additional privacy regulations and either a commissioner or ombudsperson who is responsible for the privacy legislation in that region (provinces, ). With this study, we identify and increase understanding in regards to Canadian industry’s practice and challenges in compliance with privacy regulations, both existing (PIPEDA) and emerging (Bill C27). Specifically, we address the following research questions in the context of Canada’s industry and privacy regulations.
@@
To address our research questions we employ a qualitative survey study with 22 participants. We employ thematic analysis over the collected responses to address our research questions and make the following contributions. (i) We identify insights from industry in regard to how privacy regulations impact their business and their current practices for handling sensitive data. (ii) From our respondents’ reported processes, we derive the components of the decision-making process for PETs adoption and its inclusion of both personnel inside and outside of the organizations. (iii) We determine that challenges that impede the adoption of PETs include compatibility with existing systems, feature requirements, lack of clarity for how they fulfill regulatory requirements, and the broader socio-economic system of industry.
@@
The remainder of this paper is organized as follows. We present background on PETs in Section 2 followed by the relevant related work in Section 3, and our methodology is included in Section 4. We organize our results within each research question. Section 5 is our results for RQ1, Section 6 is our results for RQ2, Section 7 is our results for RQ3, and Section 8 is our results for RQ4. We provide additional discussion on the key results and impacts of our work in Section 9 before ending with our conclusion in Section 10.
@@ 2. Background
Privacy-enhancing technologies (PETs) are technical solutions that aim at protecting privacy, a complex and evolving concept. These technologies address diverse aspects of privacy, as illustrated by Heurix et al. (HEURIX20151, ) in their 2015 taxonomy study. Their framework categorizes PETs across seven dimensions, including scenarios (e.g., untrusted client/server), aspect (identity, content, behavior), aim (indistinguishability, unlinkability, deniability, confidentiality), foundations (security model, cryptography), data (stored, transmitted, processed), trusted-third party, and reversibility. Their evaluation covers a wide range of PETs, from 1-n oblivious transfer (OT) (10.5555/36664.36681, ) in the 1980s to anonymization techniques (kanonymity02, ; ldiversity07, ) developed in the 2000s. While many of these PETs have seen practical adoption, their effectiveness can be challenged by emerging attacks and evolving system environments. For instance, k-anonymity (kanonymity02, ), initially considered a promising technique for hiding personal identity by masking or generalizing (quasi)-identifiable information, was then shown to be vulnerable to new attacks (ldiversity07, ), prompting the development of stronger PETs like differential privacy (DP) (Dwork2006DP, ). Similarly, while access control (AC) has been widely adopted by industry for decades, studies have found that the deployed AC mechanisms are not necessarily compatible with privacy regulations like GDPR (gdprac22, ), and there are challenges in unifying different AC models and techniques for enforcement purposes, particularly in federated environments (ac18, ).
This work considers three types of PETs based on their industrial deployment and research maturity. The first type includes PETs that are widely adopted in the industry but may not represent the state-of-the-art (SOTA). Examples include k-anonymization for sensitive data release and access control for managing sensitive databases. The second category comprises PETs with widespread deployment and relatively few identified critical vulnerabilities, such as multi-factor authentication (10.1145/3440712, ) for online login, SSL/TLS (10.1145/1298306.1298318, ) for data transmission, and AES encryption (HERON20098, ) for data storage. Finally, the third category consists of SOTA PETs, which are actively researched but not yet broadly deployed in the industry. These include differential privacy (Dwork2006DP, ), multi-party computation (mpc87, ), oblivious RAMs (oram96, ; oram19, ), homomorphic encryption (he15, ), trusted execution environments(TEEs) (tee15, ), federated learning (fl21, ), and synthetic data (JPMorgan2024synthetic, ; NIST2018, ). Our study focuses on understanding what challenges impair industry adoption of these emerging PETs.
@@
PETs experts advocate for the adoption of PETs within the industry due to their multifaceted benefits. Experts believe PETs primarily empower organizations to achieve and maintain compliance with evolving privacy regulations through technical tools for anonymization, data minimization, and secure data processing (garrido2023lessons, ; birrell2024sok, ). With increasing data breaches and privacy concerns, businesses prioritizing robust data protection gain a significant competitive advantage, thus fostering customer trust and loyalty (garrido2023lessons, ). Furthermore, PETs, like differential privacy and secure multi-party computation, facilitate secure data sharing and collaborative analysis of sensitive data with both internal and with external partners in sectors such as healthcare, finance, and research (NIST2018, ; cummings2023centering, ). Finally, technologies such as oblivious RAM and homomorphic encryption enable organizations to confidently outsource sensitive data and computations backed by strong security guarantees (oram19, ; ZHAOE201973, ).
@@ 3. Related work
We give an overview of two lines of related research that intersect with our exploratory work. These two lines investigate: i) the challenges in the evolution of privacy policies from legislation to practice and ii) the transition of PETs from research to industry applications. We also explore their intersection, with a particular focus on studies examining how PETs can support policy requirements.
Privacy policies and regulations in practice.
The introduction of the European General Data Protection Regulation (GDPR) (gdpr2016, ) in 2018 brought significant changes to the data protection landscape. Similarly, California’s Consumer Privacy Act (CCPA), enacted two years later, represented another major shift in privacy regulation at the state level (ccpa2018, ). These laws, GDPR and CCPA, have seen significant study as accounted by Birrel et al. in their systematization of 24 such laws and a corresponding 270 research papers considering these laws through various lenses, including legality, compliance, and business risk (birrell2024sok, ).
The challenges of working with GDPR, its implementation process, and compliance verification received significant attention since the introduction of the regulation (GDPR_TIKKINENPIRI2018134, ; GDPR_MPS18, ), and remain an active area of research (bouhoula2024automated, ; kubicek2024automating, ; klein2023general, ; reitinger2024understanding, ; ferreira2023rulekeeper, ). Marjanov et al. (marjanov2023data, ) point out the uncertainties in compliance with GDPR even after years of being in force. They reverse engineer the fines imposed for the violations in their study to discover the characteristics of each class of violations and provide a danger-points checklist for practitioners. Regulatory uncertainty may be further exacerbated by miscommunication between experts from different fields. Horstmann et al. (Horstmann2024ThoseTA, ) conducted several interviews to investigate the communication gap between software developers and privacy experts, resulting in companies regularly being fined for violations due to noncompliance with privacy data regulations. To explore the developers’ understanding of privacy requirements, Tahaei et al. (Tahaei2022, ) analyzed numerous privacy-related accepted answers on Stack Overflow and extracted advice groups from these answers. The advice mostly focuses on the inform, hide, control, and minimize of the Hoepman’s privacy design strategies to comply with regulations. The lack of clarity can extend all the way to app policy documents, leading to non-transparency. These documents serve as the de facto agreement between developers and users and are an important basis for resolving conflicts in the event of a data breach. Yan et al. (yan2024quality, ) carried out a study on the quality of privacy policy documents in the virtual personal assistant app domain to explore the issue.
The outlined challenges undermine the effectiveness of the privacy regulations, which has been further observed by work on compliance and effectiveness of regulations (van2022setting, ; tran2024measuring, ; samarin2023lessons, ). For instance, Samarin et al. (samarin2023lessons, ) investigated the extent to which Android app developers comply with the provisions of the CCPA that require them to provide consumers with accurate privacy notices and disclose personal information that they have collected, used, or shared about consumers for a business or commercial purpose upon request. They found out that a significant percentage of apps collected information that was not disclosed.
PETs deployment challenges.
Differential privacy (DP) (Dwork2006DP, ), a promising technique that supports anonymity by protecting individuals against singling out, has been the center of attention (cummings2023centering, ; Dwork2019DPinPractice, ; garrido2023lessons, ), including from NIST which produced a set of guidelines for evaluating DP guarantees (near2023guidelines, ). Dwork et al. study deployment challenges by gathering DP experts’ opinions in 2019 (Dwork2019DPinPractice, ) and distill them into four main categories. Their findings were as well mirrored in the results of a study by Garrido et al. in 2023 (garrido2023lessons, ) on industry participants who were not experts in differential privacy: “(i) Part of the challenges in deploying DP were design based. DP tools still require engineering effort to be easily deployable in organizations. (ii) DP deployment complexity is also institutionally based. A common theme of the interviewed companies was their intricate networks of stakeholders and processes, which hamper goal alignment and technology deployments. (iii) There was no consistency in DP approaches across institutions, indicating a need for shared learning. (iv) Transparency and testable statements about privacy can benefit companies in the regulatory landscape”.
@@
Several studies evaluate or enhance the usability, deployability, and communicative ease of privacy-preserving applications for real-world scenarios (nanayakkara2024consider, ; song2024inherently, ; cummings2021need, ; nanayakkara2023chances, ; franzen2022private, ; karegar2022exploring, ; ngong2024evaluating, ; kacsmar2023comprehension, ). Qin et al. (Qin2019FromUsability, ) yield insights about the interplay between usability and security in multi-party computation (MPC) applications. Agrawal et al. (Agrawal2021Exploring, ) focus on the journey of a number of privacy-preserving computation (PPC) techniques – namely, homomorphic encryption (HE), secure MPC, and differential privacy – from research into production code. They focus on specific application contexts; usability of the libraries and tools from a non-specialist developer’s perspective; and the explanation and governance challenges associated with these techniques. Since PETs development relies heavily on cryptography in several examples, we also include the challenges in cryptography’s journey from research paper to products. Fischer et al. (Fischer2024Challenges, ) identify these challenges as: misunderstandings and miscommunication among stakeholders, unclear delineation of responsibilities, misaligned or conflicting incentives, and usability.
@@
The introduction of privacy regulations has made evident changes in PETs research papers. Some discuss how their proposed technology aligns with privacy regulations. For example, synthetic data generation technologies are perceived to provide a practical replacement for the original sensitive data (NIST2018, ; JPMorgan2024synthetic, ). Others draw inspiration from the tension between data privacy policies and socially beneficial analytics to develop usable, privacy-preserving systems – Synq (Espiritu2024Synq, ) for instance, exemplifies this in computation over encrypted data. Additionally, some PET research pays dedicated attention to the technology’s relationship with regulations. Walsh et al. (Walsh2022Multi-Regulation, ) examine privacy laws and regulations that limit disclosure of personal data, and explore whether and how these restrictions apply when participants use cryptographically secure multi-party computation (MPC). Another body of work focuses on assessing the privacy promises of PETs (stadler2022synthetic, ; Giomi2023AUnified, ). Giomi et al. introduce Anonymeter, a framework for quantifying privacy risk in synthetic data (Giomi2023AUnified, ), equipped with attack-based evaluations for the singling out, linkability, and inference risks, which are the three key indicators of factual anonymization according to data protection regulations, such as GDPR.

### v2

@@ Abstract.
Privacy is an instance of a social norm formed through legal, technical, and cultural dimensions. Institutions such as regulators, industry, and researchers act as societal agents that both influence and respond to evolving norms. Attempts to promote privacy are often ineffective unless they account for this complexity and the dynamic interactions among these actors. Privacy enhancing technologies (PETs) are technical solutions for privacy issues that enable collaborative data analysis, allowing for the development of solutions that benefit society, all while ensuring the privacy of individuals whose data is being used. However, despite increased privacy challenges and a corresponding increase in new regulations being proposed by governments across the globe, a low adoption rate of PETs persists. In this work, we investigate the factors influencing industry’s decision-making processes around PETs adoption as well as the extent to which privacy regulations inspire such adoption. We conducted a qualitative survey study with 22 industry participants from across Canada to investigate how businesses in Canada make decisions to adopt novel technologies and how new privacy regulations impact their business processes. Informed by the results of our analysis, we make recommendations for industry, researchers, and policymakers on how to support what each of them seeks from the other when attempting to improve digital privacy protections. By advancing our understanding of what challenges the industry faces, we increase the effectiveness of future privacy research that aims to help overcome these issues.
Privacy, Industry Adoption, Regulations, Privacy-Enhancing Technologies
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/ccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
Societal demand for privacy is influencing law-makers, resulting in ongoing introductions of new privacy regulations globally (of Caliornia, 2018; Union, 2016; of Canada, 2022; Edwards, 2021). Most Canadian companies are moderately aware of their responsibilities under Canada’s privacy laws and have taken steps to ensure they comply with these laws according to the most recent survey (2023-2024) of Canadian businesses (Phoenix Strategic Perspectives Inc., 2024). However, while there has been a lot of effort in privacy enhancing technologies (PETs) research to develop deployable privacy-enhancing tools that aid organizations in their privacy preservation approaches, it is unclear whether additional regulations will motivate increased usage of these tools by industry. The divide between research and practice is a prolific one, and it is known that technical developments and the social components of their use in practice can have a great divide (Ackerman, 2000).
The implications of social contexts for knowledge sharing is an unavoidable first hurdle to bridging the divide before the adoption of novel technologies (Ackerman et al., 2003, 2013). In high risk domains, such as emergency rooms, it has been established that better systems emerge when consideration is given to how systems get to those who need them and how to meet those users needs (Normark and Randall, 2005). Thus, researchers need to understand how industry approaches expertise collection if we are going to improve and resolve the gaps and issues originating from different expectations on expertise communication and gathering strategies (Pipek et al., 2012; Halverson et al., 2004).
We investigate the relationship of technical developments of PETs to the social practice of deployment of PETs in conjunction with how emergent policies and regulations are influencing the adoption or avoidance of such technologies. In doing so we observe PETs to be a notable example of the interconnected nature of policy, practice and design for systems aimed at social-technical problems like digital privacy (Jackson et al., 2014) – which is what PETs tries to protect. Specifically, we address the following research questions in the context of Canada’s industry and privacy regulations.
@@
adopt a pet arxiv
We highlight the social challenges expressed by industry as well as propose ways to overcome these issues that contribute to such divides. We focus on the identification of existing or perceived impediments to adopting these technologies and complying with regulations to better understand how the industry-research relationship can be improved. To address our research questions we employ a qualitative survey study with 22 participants. We employ thematic analysis over the collected responses to address our research questions and make the following contributions. (i) We identify insights from industry in regard to how privacy regulations impact their business and their current practices for handling sensitive data. (ii) From our respondents’ reported processes, we derive the components of the decision-making process for PETs adoption and its inclusion of both personnel inside and outside of the organizations. (iii) We determine that challenges that impede the adoption of PETs include compatibility with existing systems, feature requirements, lack of clarity for how they fulfill regulatory requirements, and the broader socio-economic system of industry.
@@
The remainder of this paper is organized as follows. We present background on PETs and Canada in Section 2 followed by the relevant related work in Section 3, and our methodology is included in Section 4. We organize our results within each research question. Section 5 is our results for RQ1, Section 6 is our results for RQ2, Section 7 is our results for RQ3, and Section 8 is our results for RQ4. We provide additional discussion on the key results and impacts of our work in Section 9 before ending with our conclusion in Section 10.
@@ 2. Background
Privacy-enhancing technologies (PETs) are technical solutions that aim at protecting privacy, a complex and evolving concept. These technologies address diverse aspects of privacy, as illustrated by Heurix et al. (Heurix et al., 2015) in their 2015 taxonomy study. Their framework categorizes PETs across seven dimensions, including scenarios (e.g., untrusted client/server), aspect (identity, content, behavior), aim (indistinguishability, unlinkability, deniability, confidentiality), foundations (security model, cryptography), data (stored, transmitted, processed), trusted-third party, and reversibility. Their evaluation covers a wide range of PETs, from 1-n oblivious transfer (OT) (Brassard et al., 1987) in the 1980s to anonymization techniques (Sweeney, 2002; Machanavajjhala et al., 2007) developed in the 2000s. While many of these PETs have seen practical adoption, their effectiveness can be challenged by emerging attacks and evolving system environments. For instance, k-anonymity (Sweeney, 2002), initially considered a promising technique for hiding personal identity by masking or generalizing (quasi)-identifiable information, was then shown to be vulnerable to new attacks (Machanavajjhala et al., 2007), prompting the development of stronger PETs like differential privacy (DP) (Dwork, 2006). Similarly, while access control (AC) as a PET (Borking et al., 2003) has been widely adopted by industry for decades, studies have found that the deployed AC mechanisms are not necessarily compatible with privacy regulations like GDPR (Daoudagh and Marchetti, 2022), and there are challenges in unifying different AC models and techniques for enforcement purposes, particularly in federated environments (Colombo and Ferrari, 2018).
This work considers three types of PETs based on their industrial deployment and research maturity. The first type includes PETs that are widely adopted in the industry but may not represent the state-of-the-art (SOTA). Examples include k-anonymization for sensitive data release and access control for managing sensitive databases. The second category comprises PETs with widespread deployment and relatively few identified critical vulnerabilities, such as multi-factor authentication (Jacomme and Kremer, 2021) for online login, SSL/TLS (Lee et al., 2007) for data transmission, and AES encryption (Heron, 2009) for data storage. Finally, the third category consists of SOTA PETs, which are actively researched but not yet broadly deployed in the industry. These include differential privacy (Dwork, 2006), multi-party computation (Goldreich et al., 1987), oblivious RAMs (Goldreich and Ostrovsky, 1996; Tople et al., 2019), homomorphic encryption (Armknecht et al., 2015), trusted execution environments (TEEs) (Sabt et al., 2015), federated learning (Kairouz et al., 2021), and synthetic data (Potluru et al., 2024; National Institute of Standards and Technology, 2018). Our study focuses on understanding what challenges impair industry adoption of these emerging PETs.
@@
PETs experts advocate for the adoption of PETs within the industry due to their multifaceted benefits. They believe PETs primarily empower organizations to achieve and maintain data protection and compliance with evolving privacy regulations through technical tools for anonymization, data minimization, and secure data processing (Garrido et al., 2023; Birrell et al., 2024). With increasing data breaches and privacy concerns, businesses prioritizing robust data protection gain a significant competitive advantage, thus fostering customer trust and loyalty (Garrido et al., 2023). Furthermore, PETs, like differential privacy and secure multi-party computation, facilitate secure data sharing and collaborative analysis of sensitive data with both internal and with external partners in sectors such as healthcare, finance, and research (National Institute of Standards and Technology, 2018; Cummings and Sarathy, 2023). Finally, technologies such as oblivious RAM and homomorphic encryption enable organizations to confidently outsource sensitive data and computations backed by strong security guarantees (Tople et al., 2019; Zhao E and Geng, 2019).
Canadian privacy regulations.
Personal Information Protection and Electronic Documents Act (PIPEDA) (of Canada, 2000) is Canada’s federal privacy law for private-sector organizations. It sets out the ground rules for how businesses must handle personal information in the course of their commercial activity. The Canadian government proposed Bill C-27 (of Canada, 2022) in 2022 to enact the Consumer Privacy Protection Act, the Personal Information and Data Protection Tribunal Act and the Artificial Intelligence and Data Act (AIDA). While, the federal privacy law reforms and artificial intelligence regulation contained in the bill are put in an uncertain state following the prorogation of Parliament in January 2025, their potential impact on Canada’s industry cannot be ignored (LaCasse, 2025). Further, each of Canada’s provinces and territories has their own additional privacy regulations and either a commissioner or ombudsperson who is responsible for the privacy legislation in that region (of the Privacy Commisioner of Canada, 2024). Finally, we note that in Canada, the privacy regulations apply to all sectors and are not separated out by areas such as health having their own regulations. Rather, all sectors must protect that which is deemed to be potentially privacy-harmful regardless of whether the sector that initiated the data collection is a sector considered to be high-risk or privacy critical.
@@ 3. Related work
In our work, we center the participants’ views and experiences, rather than centering our own preconceived notions or those of the broader research community, to allow us to highlight the differences of interpretations of PETs terminologies (e.g., PETs itself, anonymization, best practices for privacy, etc.) between industry and researchers. Prior research mainly focuses on evaluating comprehension or expectations for a particular PET, such as differential privacy (DP), multiparty computation (MPC), or homomorphic encryption (HE). We provide an overview of related research that seeks to motivate and facilitate the deployment of PETs by examining their interaction with key actors and influencing factors, such as industry needs and regulatory requirements.
PETs deployment challenges and facilitators.
@@
Numerous studies examine the challenges involved in deploying specific PETs within industry settings. As a promising technique that supports anonymity, Differential privacy (DP) (Dwork, 2006) has been the center of attention (Cummings and Sarathy, 2023; Dwork et al., 2019; Garrido et al., 2023). Notably, NIST produced a set of guidelines for evaluating DP guarantees (Near et al., 2023). Dwork et al. study deployment challenges by gathering DP experts’ opinions in 2019 (Dwork et al., 2019) and distill them into four main categories. Their findings were as well mirrored in the results of a study by Garrido et al. in 2023 (Garrido et al., 2023) on industry participants who were not experts in DP. Several research evaluate or enhance the usability, deployability, and communicative ease of PETs for real-world scenarios. For DP, they cover a wide range, including: user expectations of the DP mechanism  (Cummings et al., 2021) and their mental models of DP open-source libraries (Song et al., 2024), the usability of DP tools (Ngong et al., 2024), the need for explanation of the gained privacy (Nanayakkara et al., 2023) in user-suitable metaphors (Karegar et al., 2022) or in risk communication format (Franzen et al., 2022), and considerations for policy makers (Nanayakkara and Hullman, 2024). The studies of user experience extends to other forms of private computation, such as multi-party computation and private query execution as well (Kacsmar et al., 2023). Qin et al. (Qin et al., 2019) yield insights about the interplay between usability and security in multi-party computation (MPC) applications. Agrawal et al. (Agrawal et al., 2021) focus on the journey of a number of privacy-preserving computation (PPC) techniques – namely, homomorphic encryption (HE), secure MPC, and DP – from research into production code. They focus on specific application contexts; usability of the libraries and tools from a non-specialist developer’s perspective; and the explanation and governance challenges associated with these techniques. Since PETs development relies heavily on cryptography in several examples, we also include the challenges in cryptography’s journey from research paper to products. Fischer et al. (Fischer et al., 2024) identify these challenges as: misunderstandings and miscommunication among stakeholders, unclear delineation of responsibilities, misaligned or conflicting incentives, and usability.
In contrast to the prior research, our work not only includes the technical motivators to PETs adoption, but also the strategic influences – factors that determine whether PETs succeed or fail in real-world cost-benefit evaluations. We explore how industry perceives PETs, what drives their interest, and the obstacles that hinder adoption.
@@
The introduction of privacy regulations has made evident changes in PETs research papers. Some discuss how their proposed technology aligns with privacy regulations. For example, synthetic data generation technologies are perceived to provide a practical replacement for the original sensitive data (National Institute of Standards and Technology, 2018; Potluru et al., 2024). Others draw inspiration from the tension between data privacy policies and socially beneficial analytics to develop usable, privacy-preserving systems – Synq (Espiritu et al., 2024) for instance, exemplifies this in computation over encrypted data. Additionally, some PETs research gives dedicated attention to the technology’s relationship with regulations. Walsh et al. (Walsh et al., 2022) examine privacy laws and regulations that limit disclosure of personal data, and explore whether and how these restrictions apply when participants use cryptographically secure multi-party computation (MPC). Another body of work focuses on assessing the privacy promises of PETs (Stadler et al., 2022; Giomi et al., 2022). Giomi et al. introduce Anonymeter, a framework for quantifying privacy risk in synthetic data (Giomi et al., 2022), equipped with attack-based evaluations for the singling out, linkability, and inference risks, which are the three key indicators of factual anonymization according to data protection regulations, such as GDPR.
As regulations increasingly mandate the deployment of privacy-preserving mechanisms – and with the emergence of Canada’s new privacy regulation (C-27) serving as a key inspiration to this work – we included a focused investigation into how regulations influence PETs adoption.